Operation Summary for February 16, 2011 - February 18, 2011 |
|
Wednesday Feb. 16 The day shift began with the D-Zero (D0) experiment in access, with the Tevatron (TEV) off and in access, with Operations monitoring the Recycler stash (161.4E10) and the Antiproton Source (Pbar) stack (170.7E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with Electron Cooling down for repairs, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) also in access. During the day shift, operators prepared F sector for access for vacuum work. EE Support personnel replaced a bad NuMI magnet power supply. Operations reported that a NuMI water skid tripped off and required two hours to recover. Operations reported that D0 and CDF came out of access at 3:10 PM. The TEV checkout lasted through to the end of the shift. During the evening shift, operators began the TEV turn on. Operation put the TEV into HEP shot setup at 7:30 PM and established store 8507 at 8:24 PM with an average initial luminosity (AIL) of 124.9E30. EE Support personnel completed their repairs to the NuMI magnet power supply at 5 PM. Operations reported an emittance blow up in the Pbar Accumulator that caused a 23 mA loss of antiprotons.Thursday Feb. 17 The midnight shift began with Operations monitoring store 8507 (70.5E30), the stash (153.8E10), and the Pbar stack (127.6E10), with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that Pelletron personnel had started filling the tank with insulating gas (SF6) around 1 AM, but the pump-down stalled and Pelletron personnel spent the rest of the shift looking for leaks. Operations terminated store 8507 at 6:36 AM, put the TEV into shot setup, and established store 8508 at 7:45 AM with an AIL of 123.05E30. Operations established beam to a returning Meson FTBF experiment, T953, at 4:14 AM. During the day shift, Operations reported that Pelletron repairs continued. Operations placed Pbar in standby mode at 8:23 AM for stacktail equalizer studies. Stacking resumed at noon. MiniBooNE had about two hours of downtime due to a PCI interface chassis failure. At 1:58 PM, a network storm hit, which disabled beam for twenty minutes during the investigation. Operations reported that TEV personnel began an end-of-store study at 3 PM. During the evening shift, Operations terminated store 8508 at 6:30 PM, put the TEV into shot setup, and established store 8509 at 7:48 PM with an AIL of 109.77E30. Operations disabled beam to T-953 at 6 PM. Pelletron repairs continued. Operations reported multiple trips of a MI RF station (MIRF14).Friday Feb. 18 The midnight shift began with Operations monitoring store 8509 (56.03E30) and the stash (139.2E10), with Pbar stacking (92.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8509 at 4:08 AM, put the TEV into shot setup, and established store 8510 at 5:26 AM with an AIL of 100.55E30. Operations reported TEV beam position problems at the start of the store due to two issues: the D17 collimators were at the wrong positions (TEV personnel tried, but could not move the collimators) and vacuum issues at TEV sector D1. TEV personnel investigated these problems throughout the shift. Operations reported that the Linac Buncher tripped off several times during the shift. Operations established beam to T-953 at 6:55 AM. The Plans for Wednesday The Run Coordinator will allow TEV personnel to access this morning (2/18/11) to investigate the collimators. CDF and D0 will access their collision halls. The plans for the weekend will, include stacking, stashing, and storing. |
|
More Information Comments and Suggestions |
| last modified 02/18/2011 email Fermilab |
| Security, Privacy, Legal | |